OpenStudy (anonymous):

Solve the systems of equations. 2x - y = 5, 3x + 2y = 4. A. (3,1) b. (0,2) c. (2, -1) d/ (9/7, 17/7) I got C. but i'm not to sure about how to do this!

OpenStudy (mathmate):

If you think you've got C and not sure, you can substitute into the equations and see if they work: C. (2,-1) 2(2)-(-1)=4+1=5 OK 3(2)+2(-1)=6-2=4 OK So your answer is correct.